# Can LLMs Obfuscate Code? A Systematic Analysis of Large Language Models into Assembly Code Obfuscation

### v2

@@ Abstract
Malware authors often employ code obfuscations to make their malware harder to detect. Existing tools for generating obfuscated code often require access to the original source code (e.g., C++ or Java), and adding new obfuscations is a non-trivial, labor-intensive process. In this study, we ask the following question: Can Large Language Models (LLMs) potentially generate a new obfuscated assembly code? If so, this poses a risk to anti-virus engines and potentially increases the flexibility of attackers to create new obfuscation patterns. We answer this in the affirmative by developing the MetamorphASM benchmark comprising MetamorphASM Dataset (MAD) along with three code obfuscation techniques: dead code, register substitution, and control flow change. The MetamorphASM systematically evaluates the ability of LLMs to generate and analyze obfuscated code using MAD, which contains 328,200 obfuscated assembly code samples. We release this dataset and analyze the success rate of various LLMs (e.g., GPT-3.5/4, GPT-4o-mini, Starcoder, CodeGemma, CodeLlama, CodeT5, and LLaMA 3.1) in generating obfuscated assembly code. The evaluation was performed using established information-theoretic metrics and manual human review to ensure correctness and provide the foundation for researchers to study and develop remediations to this risk. The source code can be found at the following GitHub link: https://github.com/mohammadi-ali/MetamorphASM.
@@ Introduction
In dead code insertion, malware inserts sections of code that are irrelevant to the program’s normal operation. This technique can take various forms, such as adding redundant instructions, unused variables, or unreachable code branches. These additional code segments alter the structure and behavior of the malware without affecting its functionality, causing it to appear different each time it is executed. Consequently, traditional signature-based antivirus detection methods become less effective against metamorphic malware employing this technique. Listing-1 shows a snippet assembly code which we called “original code” and Listing-2 shows the original code after inserting dead code, such as ’NOP’ or ’MOV EDI, EDI’. Inserting dead code or garbage code, is one of the important method among the obfuscation techniques (Na, Choi, and Lee 2023).
@@
In this technique, the malware replaces register names used within its instructions with alternative register names. For example, if the original malware code uses the EAX register for a specific computation, the register substitution technique might replace instances of ‘EAX’ with ‘EBX’ or another available register. Although the functionality of the code remains the same, altering the register names changes the code structure from its original form. Listing 3 illustrates the register substitution technique (Balakrishnan and Schulze 2005).
@@
We conducted the benchmarking of MetamorphASM utilizing a diverse array of large language models (LLMs), which we categorized into open source (o) and proprietary (p) types, further divided into a mixture of experts (MoE) and non-mixture of experts (n-MoE) models. Access to proprietary LLMs was facilitated through APIs, which are computationally demanding and financially expensive. Therefore, we selected 15,000 assembly code samples - 5,000 per obfuscation mechanism - to ensure a fair and reasonable comparison - from our extensive repository of 300,000 examples. For our proprietary LLM evaluations, we included the GPT-4o-mini (p) (Achiam et al. 2023), an MoE model adept at handling complex assembly code patterns (OpenAI 2024). Competing against GPT-4o-mini, we employed the open-source DeepSeekCoder-v2 (o; 236B; MoE model)(Zhu et al. 2024) and Codestral (MistralAI 2024) (o; 22B; n-MoE model), both of which have undergone extensive training on assembly codes and are proficient in generating such codes. Considering their established effectiveness in advanced coding tasks, CodeLLAMA (o; 34B) (Roziere et al. 2023) and LLAMA 3.1(MetaAI 2024) (p; 405B; n-MoE) were also included in our benchmarking process. Lastly, CodeGemma (o; 7B)(Team 2024) and the trainable CodeT5 (o; 1.2B) (Wang et al. 2021) were considered models suitable for conducting future white box studies into obfuscation using LLMs. Except for CodeT5, all the LLMs were subjected to examination using zero-shot prompting (see Tables 1, 2, 3 for zero-shot prompting template) and in-context learning, with test setups including 1, 3, 5, 10, and 15 shots (see Table 4 for few shot prompting template).
Zero Shot Dead Code Prompt
Prompt: Garbage (or Dead) assembly code refers to sections of assembly code that are functionally irrelevant and do not contribute to the intended functionality of the program. These instructions are essentially “dead” because they have no impact on the program’s behavior. Examples include NOP (No Operation) instructions, MOV EAX, EAX, or ADD EAX, 0. Given the following original assembly code, randomly insert four or five garbage or dead instructions into the original code, ensuring that the overall structure and functionality of the original code are not altered. Just print the output code without any comments.
Original Assembly Code:
PUSH EDI
POP EDI
MOV EAX …
Response:
MOVZX EAX, AL
NEG EAX
ADD EAX, 0 …
Table 1: Zero Shot Dead Code Prompt Structure.
Zero Shot Register Substitution Prompt
Prompt: Register substitution in assembly code obfuscation refers to the technique of replacing one register with another throughout a section of code. The main goal is to maintain the original functionality of the code while using different registers to obscure the operations being performed. Given the following original assembly code, determine which registers can be safely substituted without altering the program’s behavior. This usually involves registers that are not used simultaneously or whose contents are not needed at the same point in the code. Substitute one register for another in the instructions. Just print the output code.
Original Assembly Code:
PUSH EDI
POP EDI
MOV EAX …
Response:
MOVZX EAX, AL
NEG EAX
ADD EAX, 0 …
Table 2: Zero Shot Register Substitution Prompt Structure.
@@
Functional correctness is an alternative method but it is impossible at this scale. Prior works use SMT solvers to validate compiler transforms are NP-Complete (Yang et al. 2024) and only guarantee soundness (i.e., no false equivalents) but not completeness (i.e., no missed equivalences), and tools used in practice expect some lifting from ASM to a higher representation (i Montolio 2023). For this reason, we use Delta Entropy, which was proposed/evaluated by Yang et al. (2024) , who found it an effective means of large-scale evaluation, as too dramatic a change in score is a strong indicator that the code has changed too significantly by either whole-re-writing (large change) or no edits (i.e., no change in score). There is a middle ground of plausibly valid transformations, and our use of manual expert evaluation over two months remediates this final uncertainty.
@@ Discussion:
Codestral stands out among specialized large language models for its effective performance in dead code and control flow change tasks, with cosine similarity values ranging from 0.90 to 0.98 (see Tables 5, 7). However, it struggles with register substitution, indicating difficulties in modifying register names effectively Table 6. DeepSeekCoder follows with higher performance, reflected in its elevated cosine similarity scores. These suggest that it accurately replicates the original assembly code, hinting at its specialized training in assembly language, making it a proficient obfuscator across all techniques. In contrast, CodeGemma and CodeLLAMA show inadequate results in the three obfuscation techniques, primarily due to their training in high-level programming languages like C/C++/C#, Java, Rust, and Python rather than assembly. This leads to significant inaccuracies and irrelevant outputs. StarCoder, while capable of generating assembly code, demonstrates a high variability in entropy, suggesting it understands assembly but fails to consistently obfuscate at this level. Similarly, control flow change is the most effective obfuscation technique across specialized LLMs, followed by dead code insertion. Register substitution appears weaker, with a higher susceptibility to de-obfuscation. CodeT5, despite being fine-tuned, produces high cosine similarity in register substitution, around 0.98, indicating a strong resemblance to the original code. Yet, its low entropy suggests minimal to no actual obfuscation, often merely reproducing the original code as obfuscated. Human Evaluation: We designed three criteria to assess the effectiveness of obfuscation techniques in various code-generating language models, each rated on a scale from 1 to 8. The criteria are: (a) ranking the eight outputs based on the insertion of ineffective code, (b) ranking based on the substitution of registers, and (c) ranking based on the rearrangement of code sequences. We chose 200 random assembly code samples for this evaluation, conducted by three experts specializing in malware analysis. The results are in Table 8, where a lower score signifies higher-quality obfuscation.

### v3

@@ Abstract
Malware authors often employ code obfuscations to make their malware harder to detect. Existing tools for generating obfuscated code often require access to the original source code (e.g., C++ or Java), and adding new obfuscations is a non-trivial, labor-intensive process. In this study, we ask the following question: Can Large Language Models (LLMs) potentially generate a new obfuscated assembly code? If so, this poses a risk to anti-virus engines and potentially increases the flexibility of attackers to create new obfuscation patterns. We answer this in the affirmative by developing the MetamorphASM benchmark comprising MetamorphASM Dataset (MAD) along with three code obfuscation techniques: dead code, register substitution, and control flow change. The MetamorphASM systematically evaluates the ability of LLMs to generate and analyze obfuscated code using MAD, which contains 328,200 obfuscated assembly code samples. We release this dataset and analyze the success rate of various LLMs (e.g., GPT-3.5/4, GPT-4o-mini, Starcoder, CodeGemma, CodeLlama, CodeT5, and LLaMA 3.1) in generating obfuscated assembly code. The evaluation was performed using established information-theoretic metrics and manual human review to ensure correctness and provide the foundation for researchers to study and develop remediations to this risk.
Code and Datasethttps://github.com/mohammadi-ali/MetamorphASM
@@ Introduction
In dead code insertion, malware inserts sections of code that are irrelevant to the program’s normal operation. This technique can take various forms, such as adding redundant instructions, unused variables, or unreachable code branches. These additional code segments alter the structure and behavior of the malware without affecting its functionality, causing it to appear different each time it is executed. Consequently, traditional signature-based antivirus detection methods become less effective against metamorphic malware employing this technique. Listing-1 shows a snippet assembly code which we called “original code” and Listing-2 shows the original code after inserting dead code, such as “NOP” or “MOV EDI, EDI”. Inserting dead code or garbage code, is one of the important method among the obfuscation techniques (Na, Choi, and Lee 2023).
@@
In this technique, the malware replaces register names used within its instructions with alternative register names. For example, if the original malware code uses the “EAX” register for a specific computation, the register substitution technique might replace instances of “EAX” with “EBX” or another available register. Although the functionality of the code remains the same, altering the register names changes the code structure from its original form. Listing 3 illustrates the register substitution technique (Balakrishnan and Schulze 2005).
@@
We conducted the benchmarking of MetamorphASM utilizing a diverse array of large language models (LLMs), which we categorized into open source (o) and proprietary (p) types, further divided into a mixture of experts (MoE) and non-mixture of experts (n-MoE) models. Access to proprietary LLMs was facilitated through APIs, which are computationally demanding and financially expensive. Therefore, we selected 15,000 assembly code samples - 5,000 per obfuscation mechanism - to ensure a fair and reasonable comparison - from our extensive repository of 300,000 examples. For our proprietary LLM evaluations, we included the GPT-4o-mini (p) (Achiam et al. 2023), an MoE model adept at handling complex assembly code patterns (OpenAI 2024). Competing against GPT-4o-mini, we employed the open-source DeepSeekCoder-v2 (o; 236B; MoE model)(Zhu et al. 2024) and Codestral (MistralAI 2024) (o; 22B; n-MoE model), both of which have undergone extensive training on assembly codes and are proficient in generating such codes. Considering their established effectiveness in advanced coding tasks, CodeLLAMA (o; 34B) (Roziere et al. 2023) and LLAMA 3.1(MetaAI 2024) (p; 405B; n-MoE) were also included in our benchmarking process. Lastly, CodeGemma (o; 7B)(Team 2024) and the trainable CodeT5 (o; 1.2B) (Wang et al. 2021) were considered models suitable for conducting future white box studies into obfuscation using LLMs. Except for CodeT5, all the LLMs were subjected to examination using zero-shot prompting (refer to Table 1 for the zero-shot control flow change prompting template. The prompt templates for Dead Code and register substitutions can be found in the supplementary code materials.) and in-context learning, with test setups including 1, 3, 5, 10, and 15 shots (see Table 2 for few shot prompting template).
@@
Functional correctness is an alternative method but it is impossible at this scale. Prior works use SMT solvers to validate compiler transforms are NP-Complete (Yang et al. 2024) and only guarantee soundness (i.e., no false equivalents) but not completeness (i.e., no missed equivalences), and tools used in practice expect some lifting from ASM to a higher representation (Montolio 2023). For this reason, we use Delta Entropy, which was proposed/evaluated by Yang et al. (2024) , who found it an effective means of large-scale evaluation, as too dramatic a change in score is a strong indicator that the code has changed too significantly by either whole-re-writing (large change) or no edits (i.e., no change in score). There is a middle ground of plausibly valid transformations, and our use of manual expert evaluation over two months remediates this final uncertainty.
@@ Discussion:
Codestral stands out among specialized large language models for its effective performance in dead code and control flow change tasks, with cosine similarity values ranging from 0.90 to 0.98 (see Tables 3, 5). However, it struggles with register substitution, indicating difficulties in modifying register names effectively Table 4. DeepSeekCoder follows with higher performance, reflected in its elevated cosine similarity scores. These suggest that it accurately replicates the original assembly code, hinting at its specialized training in assembly language, making it a proficient obfuscator across all techniques. In contrast, CodeGemma and CodeLLAMA show inadequate results in the three obfuscation techniques, primarily due to their training in high-level programming languages like C/C++/C#, Java, Rust, and Python rather than assembly. This leads to significant inaccuracies and irrelevant outputs. StarCoder, while capable of generating assembly code, demonstrates a high variability in entropy, suggesting it understands assembly but fails to consistently obfuscate at this level. Similarly, control flow change is the most effective obfuscation technique across specialized LLMs, followed by dead code insertion. Register substitution appears weaker, with a higher susceptibility to de-obfuscation. CodeT5, despite being fine-tuned, produces high cosine similarity in register substitution, around 0.98, indicating a strong resemblance to the original code. Yet, its low entropy suggests minimal to no actual obfuscation, often merely reproducing the original code as obfuscated. Human Evaluation: We designed three criteria to assess the effectiveness of obfuscation techniques in various code-generating language models, each rated on a scale from 1 to 8. The criteria are: (a) ranking the eight outputs based on the insertion of ineffective code, (b) ranking based on the substitution of registers, and (c) ranking based on the rearrangement of code sequences. We chose 200 random assembly code samples for this evaluation, conducted by three experts specializing in malware analysis. The results are in Table 6, where a lower score signifies higher-quality obfuscation.